The foot of a ladder is 1m from a vertical wall. If the ladder reaches 5m up the wall, find the length of the ladder, to the nearest cm
Answer by Theo(13342)

the ladder forms a right triangle with the wall.

the vertical length up the wall is one leg.

the horizontal length from the wall to the ladder is the other leg.

the length of the ladder is the hypotenuse.

in a right triangle, c^2 = a^2 + b^2

c is the hypotenuse
a is one leg
b is the other leg.

c^2 = a^2 + b^2 becomes c^2 = 5^2 + 1^2 = 25 + 1 = 26

c is equal to the square root of (26).

that's the length of the ladder.

square root of 26 is approximately equal to 5.099 meters rounded to 3 decimal places.